Some dude has 200 ml of a solution that is 15% alcohol. Find how many milimeters of a 40% alcohol solution he should mix into get a final mixture of 30% alcoholic
Answer by Alan3354(69443)

Some dude has 200 ml of a solution that is 15% alcohol. Find how many milimeters of a 40% alcohol solution he should mix into get a final mixture of 30% alcoholic
==========================
200*15 + x*40 = (200+x)*30
40x + 3000 = 30x + 6000
x = 300 ml